OpenStudy (anonymous):

what is 1/4 plus 5/12

OpenStudy (anonymous):

\[\frac{ 3 }{ 12 }+\frac{ 5 }{ 12 }\]

OpenStudy (anonymous):

so you would do 5 + 3, and get \[\frac{ 8 }{ 12 }\] which you would then simplify.